# The Seifert-van Kampen Theorem via Computational Paths: A Formalized Approach to Computing Fundamental Groups

### v2

@@ Abstract.
The Seifert-van Kampen theorem computes the fundamental group of a space from the fundamental groups of its constituents. We formalize this theorem within the framework of computational paths—an approach to equality where witnesses are explicit sequences of rewrites governed by the confluent, terminating $\mathrm{LND}_{\mathrm{EQ}}$-TRS.
Our contributions are: (i) pushouts as higher-inductive types with explicit path constructors; (ii) free products and amalgamated free products as quotients of word representations; (iii) an encode-decode proof establishing $\pi_{1}(\mathrm{Pushout}(A,B,C))\simeq\pi_{1}(A)\ast_{\pi_{1}(C)}\pi_{1}(B)$; and (iv) applications to the figure-eight ($\pi_{1}(S^{1}\vee S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$), 2-sphere ($\pi_{1}(S^{2})\simeq 1$), and 3-sphere ($\pi_{1}(S^{3})\simeq 1$ with Hopf fibration context).
Recent extensions include: higher homotopy groups $\pi_{n}$ via the weak $\omega$-groupoid structure (with $\pi_{2}$ abelian via Eckmann-Hilton, and $\pi_{2}\cong 1$ in the 1-groupoid truncated setting); truncation levels connecting the framework to HoTT n-types; automated path simplification tactics; basic covering space theory with $\pi_{1}$-actions on fibers; fibration theory with long exact sequences; and Eilenberg-MacLane space characterization ($S^{1}=K(\mathbb{Z},1)$).
The framework makes coherence witnesses explicit as rewrite derivations. The development is formalized in Lean 4 with over 25,000 lines of mechanized proofs. This demonstrates that the encode-decode method for higher-inductive types becomes fully constructive when path equality is decidable via normalization.
@@ 1. Introduction
The present paper develops the Seifert-van Kampen theorem within the framework of computational paths [2, 13]—an alternative approach to equality in type theory where witnesses of equality are explicit sequences of rewrites. This approach offers several advantages:
@@
    Decidable equality on paths: The rewrite equality relation ($\sim$) is decidable via normalization, making path equality computationally tractable.
@@ 1.1. Main Contributions
    A formal definition of pushouts as higher-inductive types within the computational paths framework, including:
    • Point constructors $\mathrm{inl}:A\to\mathrm{Pushout}$ and $\mathrm{inr}:B\to\mathrm{Pushout}$
    • Path constructor $\mathrm{glue}:\prod_{c:C}\mathrm{Path}(\mathrm{inl}(f(c)),\mathrm{inr}(g(c)))$
    • Recursion and induction principles with computation rules
@@
    The construction of free products $G_{1}\ast G_{2}$ and amalgamated free products $G_{1}\ast_{H}G_{2}$ as quotients of word representations.
@@
    A complete encode-decode proof of the Seifert-van Kampen theorem:
@@
    • The 2-sphere: $\pi_{1}(S^{2})\simeq 1$ (trivial group)
    • The 3-sphere: $\pi_{1}(S^{3})\simeq 1$ with Hopf fibration context
@@
    A complete Lean 4 formalization available at [15].
@@ 1.2. Related Work
The connection between rewriting and coherence has been developed by Kraus and von Raumer [9], who showed that confluence and termination suffice to build coherent higher-dimensional structure. Their key insight is that in a confluent, terminating rewrite system, any two parallel reductions to normal form can be connected by a canonical “diamond” of rewrite paths—and this extends to higher dimensions. For our $\mathrm{LND}_{\mathrm{EQ}}$-TRS, this means that higher coherences (such as the pentagon identity for associativity and the naturality conditions for transport) are automatically satisfied once we establish that paths with the same normal form are $\sim$-equivalent. This principle is crucial for the SVK proof, where glue naturality and amalgamation compatibility require reasoning about 2-dimensional path equalities.
@@ 1.4. Structure of the Paper
Section 2 reviews computational paths, rewrite equality, and fundamental groups. Section 3 defines pushouts as higher-inductive types. Section 4 constructs free products and amalgamated free products. Section 5 proves the Seifert-van Kampen theorem using the encode-decode method. Section 6 applies the theorem to compute fundamental groups of the figure-eight, 2-sphere, and 3-sphere (with Hopf fibration context). Section 7 discusses the Lean 4 formalization. Section 8 concludes.
@@ 2.1. Computational Paths
In Lean, a path is represented as a wrapper around an equality proof:
@@
2 toEq : a = b -- the underlying equality proof
Remark 2.1 (Implementation Note).
The Lean implementation stores only the underlying equality proof, not an explicit rewrite sequence. The “rewrite sequence” interpretation is recovered via the Step and RwEq relations, which classify paths syntactically. This design choice simplifies the formalization while preserving the key property that path equality is captured by rewrite equivalence.
@@ 2.2. The Step Relation
  • Canonicalization: $p\rightsquigarrow\mathrm{ofEq}(p.\mathrm{toEq})$ — any path can reduce to its canonical form.
Remark 2.2 (Critical Pairs).
The confluence and termination of $\mathrm{LND}_{\mathrm{EQ}}$-TRS are established in [12] via analysis of 56 critical pairs. The key insight is that the canon rule—which rewrites any path $p$ to $\mathrm{ofEq}(p.\mathrm{toEq})$—acts as a “normalization shortcut” that collapses all parallel paths to a common canonical form.
In Lean, the Step relation is defined inductively:
@@ 9 Step (trans (trans p q) r) (trans p (trans q r))
10 | canon (p : Path a b) : Step p (Path.ofEq p.toEq)
11 -- ... plus type-specific and context rules (50+ total)
@@ 2.3. Rewrite Equality
The $\mathrm{LND}_{\mathrm{EQ}}$-TRS enjoys termination and confluence [14, 12], so every path has a unique normal form. This yields three important properties:
  1. (1)
    Soundness: $p\sim q$ implies $p.\mathrm{toEq}=q.\mathrm{toEq}$ (rewriting preserves the underlying equality).
  2. (2)
    Completeness: Any path $p$ is rewrite-equivalent to its canonical form, so $p\sim\mathrm{ofEq}(p.\mathrm{toEq})$.
  3. (3)
    Decidability: $\sim$ is decidable by comparing canonical forms: $p\sim q$ iff their canonical forms are identical.
Theorem 2.4 (Canonicity).
In Lean 4 (with proof-irrelevant Prop), for any paths $p,q:\mathrm{Path}(a,b)$, we have $p\sim q$.
Proof.
Since $\mathrm{Path}(a,b)$ stores a proof $p.\mathrm{toEq}:a=b$ in Lean’s proof-irrelevant $\mathsf{Prop}$, we have $p.\mathrm{toEq}=q.\mathrm{toEq}$ definitionally. The canon rule gives $p\sim\mathrm{ofEq}(p.\mathrm{toEq})$ and $q\sim\mathrm{ofEq}(q.\mathrm{toEq})$. Since $p.\mathrm{toEq}=q.\mathrm{toEq}$, these canonical forms are identical, so $p\sim q$ by transitivity. ∎
Remark 2.5 ($\sim$ Triviality in Lean).
In our Lean 4 implementation, since Path.toEq stores a proof in the proof-irrelevant Prop type, any two paths $p,q:\mathrm{Path}(a,b)$ satisfy $p.\mathrm{toEq}=q.\mathrm{toEq}$ definitionally. Thus $\sim$ is trivially satisfied for all parallel paths. The groupoid laws become derivable rather than requiring explicit rewrite sequences.
This differs from the original computational paths theory where paths carry explicit rewrite traces. The advantage is simpler formalization; the tradeoff is that we lose the ability to distinguish paths by their internal structure. Future work could implement paths as syntactic terms (not carrying equality proofs) to recover the intended distinction.
@@ 3.1. The Pushout Type
In Lean, pushouts are axiomatized as a HIT:
1axiom Pushout (A : Type u) (B : Type u) (C : Type u)
2 (f : C -> A) (g : C -> B) : Type u
3
4namespace Pushout
5 axiom inl : A -> Pushout A B C f g
6 axiom inr : B -> Pushout A B C f g
7 axiom glue (c : C) : Path (inl (f c)) (inr (g c))
8end Pushout
@@ 3.3. Glue Naturality
In Lean:
1axiom glue_natural {c1 c2 : C} (p : Path c1 c2) :
2 Path.trans (inlPath (Path.congrArg f p)) (glue c2) =
3 Path.trans (glue c1) (inrPath (Path.congrArg g p))
4
5axiom glue_natural_rweq {c1 c2 : C} (p : Path c1 c2) :
6 RwEq (inlPath (Path.congrArg f p))
7 (Path.trans (glue c1)
8 (Path.trans (inrPath (Path.congrArg g p))
9 (Path.symm (glue c2))))
@@ 4.1. Free Product Words
A word in the free product $G_{1}\ast G_{2}$ is an alternating sequence of elements from $G_{1}$ and $G_{2}$:
@@
3 | consLeft (x : G1) (rest : FreeProductWord G1 G2) : FreeProductWord G1 G2
4 | consRight (y : G2) (rest : FreeProductWord G1 G2) : FreeProductWord G1 G2
Example 4.2 (Figure-Eight Words).
For $\pi_{1}(S^{1}\vee S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$, words represent sequences of loops around the two circles:
  • $\mathrm{consLeft}(1,\mathrm{nil})$ — one loop around the left circle
  • $\mathrm{consRight}(2,\mathrm{nil})$ — two loops around the right circle
  • $\mathrm{consLeft}(1,\mathrm{consRight}(-1,\mathrm{nil}))$ — left loop, then reverse right loop
Words support concatenation:
1def concat : FreeProductWord G1 G2 -> FreeProductWord G1 G2
2 -> FreeProductWord G1 G2
3 | nil, w2 => w2
4 | consLeft x rest, w2 => consLeft x (concat rest w2)
5 | consRight y rest, w2 => consRight y (concat rest w2)
Lemma 4.3 (Word Concatenation Properties).
  1. (1)
    $\mathrm{concat}(\mathrm{nil},w)=w$
  2. (2)
    $\mathrm{concat}(w,\mathrm{nil})=w$
  3. (3)
    $\mathrm{concat}(\mathrm{concat}(w_{1},w_{2}),w_{3})=\mathrm{concat}(w_{1},\mathrm{concat}(w_{2},w_{3}))$
Free vs. Direct Product.
Unlike the torus $S^{1}\times S^{1}$ whose loops can “slide past each other” (yielding the abelian group $\mathbb{Z}\times\mathbb{Z}$), loops in $S^{1}\vee S^{1}$ must pass through the common basepoint, preventing commutation. The word $\mathrm{loop}_{A}\cdot\mathrm{loop}_{B}$ represents first traversing the left circle, then the right; reversing the order gives a genuinely distinct homotopy class $\mathrm{loop}_{B}\cdot\mathrm{loop}_{A}$.
@@ 4.2. Amalgamated Free Product
The amalgamation relation identifies elements from $H$ mapped into $G_{1}$ with their counterparts mapped into $G_{2}$:
1inductive AmalgRelation {G1 G2 H : Type u}
2 (i1 : H -> G1) (i2 : H -> G2) :
@@ 5 AmalgRelation i1 i2
8 | amalgRightToLeft (h : H) (pre suf : FreeProductWord G1 G2) :
9 AmalgRelation i1 i2
10 (concat pre (concat (singleRight (i2 h)) suf))
11 (concat pre (concat (singleLeft (i1 h)) suf))
@@ 9 AmalgRelation i1 i2
The amalgamated free product is the quotient of free product words by the equivalence closure of the amalgamation relation:
1def AmalgamatedFreeProduct (G1 G2 H : Type u)

### v3

@@ Abstract.
The Seifert-van Kampen theorem computes the fundamental group of a space from the fundamental groups of its constituents. We develop a modular SVK framework within the setting of computational paths—an approach to equality where witnesses are explicit sequences of rewrites governed by the $\mathrm{LND}_{\mathrm{EQ}}$-TRS.
Our contributions are: (i) pushouts as higher-inductive types with modular typeclass assumptions for computation rules; (ii) free products and amalgamated free products as quotients of word representations; (iii) an SVK equivalence schema $\pi_{1}(\mathrm{Pushout}(A,B,C))\simeq\pi_{1}(A)\ast_{\pi_{1}(C)}\pi_{1}(B)$ parametric in user-supplied encode/decode structure—the decode map and amalgamation-respecting properties are proved, while the encode map and round-trip properties are typeclass assumptions; and (iv) instantiations for classical spaces: figure-eight ($\pi_{1}(S^{1}\vee S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$), spheres ($\pi_{1}(S^{n})\simeq 1$ for $n\geq 2$), Klein bottle, real projective plane, orientable/non-orientable surfaces, and lens spaces.
The development is formalized in Lean 4 with 41,130 lines across 107 modules, using 36 kernel axioms for HIT type/constructor declarations. Pushout computation rules are not kernel axioms—they are typeclass assumptions, making dependencies explicit. The library also includes extended features (higher homotopy groups, covering spaces, fibrations, Mayer-Vietoris) documented in Section 9.
@@ 1. Introduction
The present paper develops a modular SVK framework within the setting of computational paths [2, 12]—an alternative approach to equality in type theory where witnesses of equality are explicit sequences of rewrites. This approach offers several advantages:
@@
    Syntactic path equality: The rewrite equality relation ($\sim$) is defined as the symmetric-transitive closure of Step rules. Under termination and confluence (established in prior work [14], not fully mechanized here), $\sim$ is decidable via normalization.
@@
  4. (4)
    Modular assumptions: The SVK framework uses split typeclass assumptions, allowing users to provide only what they need—from full encode/decode data to minimal Prop-level bijectivity claims.
@@ 1.1. Main Contributions
    Pushouts implemented via Lean’s quotient types (zero kernel axioms), with modular typeclass assumptions for computation rules:
    • Point constructors $\mathrm{inl}:A\to\mathrm{Pushout}$ and $\mathrm{inr}:B\to\mathrm{Pushout}$ (definitions, not axioms)
    • Path constructor $\mathrm{glue}:\prod_{c:C}\mathrm{Path}(\mathrm{inl}(f(c)),\mathrm{inr}(g(c)))$ (via Quot.sound)
    • Computation/naturality hypotheses as typeclasses (HasRecGlueRwEq, HasGlueNaturalLoopRwEq)
    • Automatic instance derivation for common cases (Axiom K, decidable equality, subsingletons)
@@
    The construction of free products $G_{1}\ast G_{2}$ and amalgamated free products $G_{1}\ast_{H}G_{2}$, including:
    • Standard amalgamation quotient
    • Full amalgamated free product with free group reduction (FullAmalgamatedFreeProduct)
    • Both data-level and Prop-level (choice-based) interfaces
@@
    An SVK equivalence schema parametric in user-supplied encode/decode structure, with split typeclass assumptions:
    • HasPushoutSVKEncodeQuot: encode map (assumed)
    • HasPushoutSVKDecodeEncode: decode-encode round-trip (assumed)
    • HasPushoutSVKEncodeDecode: encode-decode round-trip (assumed)
    • HasPushoutSVKDecodeAmalgBijective: Prop-level bijectivity alternative
    • Decode map and amalgamation-respecting property: proved
@@
    • Spheres: $\pi_{1}(S^{n})\simeq 1$ for $n\geq 2$
    • Real projective plane: $\pi_{1}(\mathbb{RP}^{2})\simeq\mathbb{Z}_{2}$
    • Orientable surfaces: $\pi_{1}(\Sigma_{g})\simeq\langle a_{1},b_{1},\ldots,a_{g},b_{g}\mid\prod_{i=1}^{g}[a_{i},b_{i}]=1\rangle$
    • Non-orientable surfaces: $\pi_{1}(N_{n})\simeq\langle a_{1},\ldots,a_{n}\mid a_{1}^{2}\cdots a_{n}^{2}=1\rangle$
    • Lens spaces: $\pi_{1}(L(p,q))\simeq\mathbb{Z}/p\mathbb{Z}$ for coprime $p,q$
    • Bouquet of $n$ circles: $\pi_{1}(\bigvee_{n}S^{1})\simeq F_{n}$ (free group)
@@
    A complete Lean 4 formalization comprising:
    • 41,130 lines of code across 107 modules
    • Only 36 kernel axioms (all for HITs)
    • Torus constructed as $S^{1}\times S^{1}$ (no axioms)
    • Opt-in axiom imports for assumption-free APIs
    • Automated path simplification tactics
@@ 1.2. Related Work
The connection between rewriting and coherence has been developed by Kraus and von Raumer [9], who showed that confluence and termination suffice to build coherent higher-dimensional structure. Their key insight is that in a confluent, terminating rewrite system, any two parallel reductions to normal form can be connected by a canonical “diamond” of rewrite paths—and this extends to higher dimensions. For our $\mathrm{LND}_{\mathrm{EQ}}$-TRS, this means that higher coherences (such as the pentagon identity for associativity and the naturality conditions for transport) are automatically satisfied once we establish that paths with the same normal form are $\sim$-equivalent.
@@ 1.4. Structure of the Paper
Section 2 reviews computational paths, rewrite equality, and fundamental groups. Section 3 defines pushouts as higher-inductive types with modular assumptions. Section 4 constructs free products and amalgamated free products. Section 5 proves the Seifert-van Kampen theorem using the encode-decode method with split assumptions. Section 6 applies the theorem extensively. Section 7 covers extended features including higher homotopy groups and covering spaces. Section 8 discusses the Lean 4 formalization. Section 10 concludes.
@@ 2.1. Computational Paths
In Lean, a path is represented as a structure storing both an explicit rewrite trace and an equality proof:
@@
2 steps : List (Step A) -- explicit list of rewrite steps
3 proof : a = b -- the derived equality proof
Remark 2.1 (Explicit Step Lists and Consistency).
The steps field is the distinguishing data of a path, not the underlying equality proof. Two paths with the same endpoints may have different step lists, and they are related by $\sim$ only if there is an explicit derivation transforming one step sequence into the other via the Step rules.
This design is critical for consistency: in Lean’s proof-irrelevant Prop, all equality proofs $p,q:a=b$ satisfy $p=q$ (proof irrelevance). If paths were identified solely by their equality proofs, all loops at a point would be equal, collapsing $\pi_{1}$ to the trivial group.
Instead, path identity depends on the explicit steps list. The fundamental group $\pi_{1}(A,a):=\mathrm{Loop}(A,a)/{\sim}$ quotients loops by rewrite equivalence of their step sequences—not by equality of their underlying proofs. The $\sim$ relation captures the computational content: two loops are identified only when one step sequence can be transformed to the other via the groupoid laws.
@@ 2.2. The Step Relation
  • Congruence closure: If $p\rightsquigarrow q$ then $p^{-1}\rightsquigarrow q^{-1}$, $p\cdot r\rightsquigarrow q\cdot r$, and $r\cdot p\rightsquigarrow r\cdot q$.
In Lean, the Step relation is defined inductively with 76 constructors:
@@ 9 Step (trans (trans p q) r) (trans p (trans q r))
10 | symm_congr : Step p q -> Step (symm p) (symm q)
11 | trans_congr_left (r : Path b c) : Step p q -> Step (trans p r) (trans q r)
12 | trans_congr_right (p : Path a b) : Step q r -> Step (trans p q) (trans p r)
13 -- ... plus type-specific rules (products, sums, transport, contexts)
@@ 2.3. Rewrite Equality
2.4. Metatheory of the $\mathrm{LND}_{\mathrm{EQ}}$-TRS
The rewrite system $\mathrm{LND}_{\mathrm{EQ}}$-TRS has been studied in prior work [13, 14, 2]. The following table summarizes the status of key metatheoretic properties in this formalization:
Property Status Location / Notes
Soundness Proved step_toEq in Step.lean: $p\sim q\Rightarrow p.\mathrm{proof}=q.\mathrm{proof}$
Confluence Typeclass HasJoinOfRw in Confluence.lean: users provide join witnesses
Local confluence Proved Critical pairs analyzed in Confluence.lean (specific cases)
Termination Not formalized Prior paper proofs [14]; not mechanized here
Decidability of $\sim$ Assumed Would follow from termination + confluence; not fully mechanized
Remark 2.3 (Non-symmetric Soundness).
Soundness says $p\sim q\Rightarrow p.\mathrm{proof}=q.\mathrm{proof}$. Critically, the converse does not hold: paths with the same equality proof are not necessarily $\sim$-related. This asymmetry is essential for non-trivial fundamental groups (see Remark 2.1).
@@ 2.5. Loop Spaces and Fundamental Groups
2.6. The Circle and Its Fundamental Group
Definition 2.8 (Circle HIT).
The circle $S^{1}$ is the higher-inductive type with:
  • A point constructor: $\mathrm{base}:S^{1}$
  • A path constructor: $\mathrm{loop}:\mathrm{Path}(\mathrm{base},\mathrm{base})$
Theorem 2.9 (Fundamental Group of Circle).
$\pi_{1}(S^{1},\mathrm{base})\simeq\mathbb{Z}$
This classical result, first formalized via encode-decode by Licata and Shulman [10], serves as the foundation for our SVK applications.
@@ 3.1. The Pushout Type
In Lean, pushouts are implemented using quotient types (no kernel axioms):
1-- Relation generating the pushout quotient
2inductive PushoutRel (A B C : Type u) (f : C -> A) (g : C -> B)
3 : Sum A B -> Sum A B -> Prop
4 | glue (c : C) : PushoutRel A B C f g (Sum.inl (f c)) (Sum.inr (g c))
5
6-- Pushout as quotient of Sum by PushoutRel
7def Pushout (A B C : Type u) (f : C -> A) (g : C -> B) : Type u :=
8 Quot (PushoutRel A B C f g)
9
10-- Constructors are definitions, not axioms
11def inl (a : A) : Pushout A B C f g := Quot.mk _ (Sum.inl a)
12def inr (b : B) : Pushout A B C f g := Quot.mk _ (Sum.inr b)
13def glue (c : C) : Path (inl (f c)) (inr (g c)) :=
14 Path.ofEq (Quot.sound (PushoutRel.glue c))
Remark 3.2 (Quot-Based Implementation).
Unlike Circle, Klein bottle, and other HITs which require kernel axioms for their path constructors, the Pushout uses Lean’s built-in Quot type. This adds zero kernel axioms to the trusted base. The glue path is constructed via Quot.sound, which witnesses that related elements become equal in the quotient.
@@ 3.2. Recursion Principle
3.3. Modular Typeclass Assumptions
A key design choice in our formalization is to make the computation and naturality rules typeclass assumptions rather than kernel axioms. This allows users to provide only what they need:
1class HasRecGlueRwEq (D : Type u) (onInl : A -> D) (onInr : B -> D)
2 (onGlue : forall c, Path (onInl (f c)) (onInr (g c))) where
3 rec_glue_rweq : forall c, RwEq (congrArg rec (glue c)) (onGlue c)
4
5class HasGlueNaturalLoopRwEq (c0 : C) where
6 glue_natural_loop : forall (p : LoopSpace C c0),
7 RwEq (inlPath (congrArg f p))
8 (trans (glue c0) (trans (inrPath (congrArg g p)) (symm (glue c0))))
3.4. Automatic Instance Derivation
The framework provides automatic instances for common cases:
1-- When both A and B satisfy Axiom K
2instance [DecidableEq A] [HasDecidableEqAxiomK A]
3 [DecidableEq B] [HasDecidableEqAxiomK B] :
4 HasGlueNaturalLoopRwEq c0 where
5 glue_natural_loop := fun p => by path_simp
6
7-- When C is a subsingleton (e.g., Unit for wedge sums)
8instance [Subsingleton C] : HasGlueNaturalLoopRwEq c0 where
9 glue_natural_loop := fun p => by simp [Subsingleton.elim]; path_rfl
@@ 3.6.2. Suspension
3.6.3. Bouquet of $n$ Circles
The bouquet of $n$ circles $\bigvee_{n}S^{1}$ generalizes the wedge sum:
1def BouquetN (n : Nat) : Type u :=
2 -- n circles wedged together at a single point
3 Pushout (Fin n -> Circle) PUnit (Fin n)
4 (fun i => Circle.circleBase)
5 (fun _ => PUnit’.unit)
@@ 4.1. Free Product Words
A word in the free product $G_{1}\ast G_{2}$ is an alternating sequence of elements:
@@
3 | consLeft (x : G1) (rest : FreeProductWord G1 G2)
4 | consRight (y : G2) (rest : FreeProductWord G1 G2)
The group operations on words are defined by concatenation and inversion with appropriate reductions.
Definition 4.2 (Word Operations).
1-- Concatenation
2def concat : FreeProductWord G1 G2 -> FreeProductWord G1 G2 ->
3 FreeProductWord G1 G2
4 | nil, w => w
5 | consLeft x rest, w => consLeft x (concat rest w)
6 | consRight y rest, w => consRight y (concat rest w)
7
8-- Inversion
9def invert : FreeProductWord G1 G2 -> FreeProductWord G1 G2
10 | nil => nil
11 | consLeft x rest => concat (invert rest) (singleLeft (-x))
12 | consRight y rest => concat (invert rest) (singleRight (-y))
@@ 4.2. Amalgamated Free Product
1inductive AmalgRelation (i1 : H -> G1) (i2 : H -> G2) :
@@ 4 AmalgRelation i1 i2
1def AmalgEquiv (i1 : H -> G1) (i2 : H -> G2) :=
2 EqvGen (AmalgRelation i1 i2)
3
4def AmalgamatedFreeProduct (G1 G2 H : Type u)